MAT 126, CALCULUS B, SECTIONS 1-4, FALL 2000 Course web page: http://www.math.sunysb.edu/^{∼}mde/126F 00/126F 00.html Changes to this syllabus are possible: they will be announced in class and, if appropriate,

on the web page. Students are responsible for announcements made in class. They should also check the web-page periodically.

MAT 126 is the second semester of the three-semester calculus sequence MAT 125, 126, 127. We will study integration. We will often need to use the theory of differentiation (MAT 125) and trigonometry.

PREREQUISITES: A grade of C or higher in MAT 124 or 125 or 131 or 141; or level 6 on the Mathematics Placement Examination. This exam will be given on several dates in September. Check the Math Undergraduate Office (P-144 Math Tower, phone 2-8250) for times and places.

LECTURES: Tu-Th 8:20–9:40, OLD ENG 143 INSTRUCTOR: Mark de Cataldo (coordinator) TEXTBOOK: Chapters 5 and 6 of Calculus, Concepts and Contexts, by James Stewart, Brooks/Cole

1997.

RECITATIONS: 01 F 09:25-10:20 LIB N4072, Wang; 02 M 09:25-10:20 LIB N4072, Wang; 03 W 04:25-05:20 pm LIB N4072, Valdez; 04 TH 05:30-06:25 pm SBS N436, St. John; 05 M 04:25 - 05:20 pm LIB N4072, Valdez; 06 TU 09:50 - 10:45 LIB N4072, vanCoevering; 07 TH 09:50 - 10:45 LIB N4072, vanCoevering; 08 TU 05:30 - 06:25 pm SBS N436, St. John

MAT 126 STAFF

Instructors Mark de Cataldo (Lec 01-04) : Office: MATH 3-115; Phone: 2-8262; e-mail: mde@math.sunysb.edu Office hours: Tu 10:00-11:00 am, MATH 3-115; Th 12:45-13:45 PHYS A127. Ian Dowker (Lec 05-08) : Office: MATH 4-105; Phone: 2-4005; e-mail: dowker@math.sunysb.edu Office hours: M 1:00-2:00 pm in MATH 4-105; W 1:00-2:00 PHYS A127. Harish Seshadri (Lec 09): Office: MAT 2-122; e-mail: harish@math.sunysb.edu Office hours: MandW 6-7pm. Teaching Assistants Anthony St. John (Rec 04/08) : e-mail: astjohn@mathlab.sunysb.edu Office hours: TandTH 3:30-5:30 PHYS A127. Rogerio Valdez (Rec 03/05) : Office: MATH 2-112; e-mail: valdez@math.sunysb.edu Office hours: M 12:00-2:00 PHYS A127; TuandTh 11:30-12:30 MATH 2-112. Craig VanCoevering (Rec 06/07) : Office: MAT 2-104; e-mail: craig@math.sunysb.edu Office hours: M 11-12 PHYS A127 and 12-1 MATH 2-104; W 12-1 PHYS A127 and 1-2 math 2-104. Hong-Yuan Wang (Rec 01/02) : Office: MAT 4-122; e-mail: wanghy@math.sunysb.edu Office hours: W 1:20-3:20 PHYS A127; W 3:20-5:20 MAT 4-122.

SCHEDULE OF EXAMS

MIDTERM 1: Monday October 16, 2000; 08:30-10:00 pm, place to be announced.

MIDTERM 2: Thursday November 16, 2000; 08:30-10:00 pm, place to be announced.

FINAL: WEDNESDAY DECEMBER 20; 2:00-4:30 pm, place to be announced.

IMPORTANT. Make sure that you will be available at these times. There will be no make-ups for missed mid-term exams. However, if you miss an exam for an acceptable reason, and provide me with an acceptable written excuse, the relevant mid-term will be ‘dropped’ (ignored) in computing your course grade. A letter stating that you were seen by a doctor or other medical personnel is NOT an acceptable document, unless it states that it was reasonable/proper for you to seek medical attention and medically necessary for you to miss the exam. (for privacy reasons this note/letter need not state anything beyond this). Incompletes will be granted only if documented circumstances beyond your control prevent you from taking the final exam.

HOMEWORK: Assigned weekly by the T.A. during recitation and also posted on the web. Due the following week during recitation. Only some of the problems will be graded, but which ones will not be announced in advance. Late homework will not be accepted. No exceptions.

QUIZZES: Quizzes will be given twice a month during recitation.

GRADE: Midterm 1 = 20%, Midterm 2 = 20%, Final= 40%, Homework = 10%, Quizzes = 10%. Maximum scores: Midterms 1 and 2: 200pts each ; each homework: 10pts (the best ten are used in the grade); each quiz: 20 pts (the best five are used in the grade); final 400pts. Total maximum: 1000pts. The numerical grade will be converted to a final letter grade only AFTER the final test has been graded. However, after each midterm an approximate letter grade will be given to you.

CALCULATORS: They are not allowed during exams. Books, notes etc. are not allowed during exams. If you need to bring a backpack etc. to an exam, you must place it, all pockets and flaps closed in plain sight of the proctors. If you need to open it (for example if you need a pen) ask the proctors first.

CONTACTING THE INSTRUCTOR AND/OR THE TEACHING ASSISTANT. E-mail is the best method for making appointments outside normal office hours. E-mail is not, however, a good way to ask math questions, as our typing abilities are very limited. After the course is over, if you have any questions about your final grade send a letter (not an e-mail) to your instructor, c/o Dept. Math, SUNY Stony Brook, Stony Brook N.Y. 11794-3651. You will receive a written reply. These matters will be dealt with in writing only; that way, we have a written record of what the student says, and what we reply.

GOLDEN RULES. Punctuality: no late arrivals, no early departures: they are disruptive. If, occasionally, you need to leave early, let me know beforehand. Silence: it is always a good rule and even more important for us since it is a big class; do show respect to other fellow students by not disturbing the class. NO CELLULAR PHONES.

TENTATIVE WEEK-BY-WEEK SYLLABUS, EXAMS AND PRACTICE PROBLEMS

WEEKLY SYLLABUS Sep 05 : 5.1; Sep 12 : 5.1 (cont.) 5.2; Sep 19 : 5.2 (cont.), Review of derivatives and 4.9; Sep 26 : 5.3;

Oct 03 : 5.4; Oct 10 : 5.5; Oct 17 : 5.5 (cont.), 5.6; Oct 24 : 5.7 (omit : pgs 411-413); Oct 31 : 5.8; Nov 07 : 5.9; Nov 14 : 6.1 6.2; Nov 21 : 6.2 (cont.); Nov 28 : 6.3 6.4; Dec 05 : 6.5 6.7 (omit : pgs. 479-482 and 6.6); Dec 12 : Review

EXAMS

Midterm I covers material in sections 5.1, 5.2, 5.3. Midterm II covers material in sections 5.4, 5.5, 5.6, 5.7. The Final covers material from the 2 Midterms and in addition material in sections 5.8, 5.9, 6.1, 6.2,

6.3, 6.4, 6.5 and section 6.7.

PRACTICE PROBLEMS Text: J. Stewart, Calculus, Concepts and Contexts Here is a list of homework problems. Some of them

will be assigned to you weekly. Note that each chapter ends with a review section: use it!

5.1 : 1-3,9-14,18; 5.2 : 1,2,4-8,11-18,21-36; 4.9 : 1-10,13-23,31-36; 5.3 : 1-4,7-30,43-55; 5.4 : 1-6,8,11-25,27-29; 5.5 : 1-32,37-52,53,54,57,61-63; 5.6 : 1-28,37-41; 5.7 : 1-24; 5.8 : 2,5-12,20-23,25; 5.9 : 1-3,5-32,41-46,60; 6.1 : 1-16,22-23,25,27,34-39; 6.2 : 1-7,9-12,17-29; 6.3 : 1,2,7-9,11,18,19; 6.4 : 1-6,9-14; 6.5 : 1-12,14,17-22; 6.7 : 1-9.

SPECIAL NEEDS. If you have a physical, psychiatric, medical, or learning disability that may affect your ability to carry out the assigned course work, please contact the office of Disabled Student Services (DSS), Humanities Building, room 133, telephone 632-6748/TDD. DSS will review your concerns and determine, with you, what accomodations are necessary and appropriate. All information and documentation of disability is confidential.